How much 15% acid solution should be added to 40 ml of a 6% solution to get a 7% acid solution?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

How much 15% acid solution should be added to 40 ml of a 6% solution to get a 7% acid solution?
---------------
Equation:
acid + acid = acid
0.06*40 + 0.15x = 0.07(40+x)
----
Multiply thru by 100 to get:
6*40 + 15x = 7*40 + 7x
8x = 40
x = 5 ml (amt of 15% solution needed for the mix)
===================
Cheers,
Stan H.

Answer by nerdybill(7384)

How much 15% acid solution should be added to 40 ml of a 6% solution to get a 7% acid solution?
.
Let x = amount of 15% acid solution added
then
.15x + .06(40) = .07(40+x)
.15x + 2.4 = 2.8 + .07x
.08x + 2.4 = 2.8
.08x = 0.4
x = 0.4/.08
x = 5 ml